# MonoPatchNeRF: Improving Neural Radiance Fields with Patch-based Monocular Guidance

### v1

@@ 3.3 Patch-based Photometric Consistency over Virtual Views
NeRF models tend to be under-constrained when the input images are sparse, which often leads to artifacts in novel view renderings despite a low training view loss. To mitigate this, we employ patch-based photometric consistency regularization.
As visualized in Fig 3, for each patch $P$ with associated colors $\mathbf{\hat{c}}_{p}$}, pose $\mathbf{o}_{p}$, directions $\mathbf{v}_{p}$} and rendered depth {$d_{p}$} in the training set, we randomly sample a corresponding virtual patch $P^{\ast}$ from a nearby virtual view. The correspondence is determined by unprojecting the pixel from the training view into 3D points, denoted as $\mathbf{X}_{p}=\mathbf{o}_{p}+d_{p}\mathbf{v}_{p}$, and then back-projecting $\mathbf{X}_{p}$ into the pose $\mathbf{o}_{p^{\ast}}$ of $P^{\ast}$ for the corresponding direction $\mathbf{v}_{p^{\ast}}=(\mathbf{X}_{p}-\mathbf{o}_{p^{\ast}})/\|(\mathbf{X}_{p}-\mathbf{o}_{p^{\ast}})\|$. Given $\mathbf{o}_{p^{\ast}}$ and $\mathbf{v}_{p^{\ast}}$, we predict the color $\mathbf{c}_{p^{\ast}}$ and depth ${d}_{p^{\ast}}$, and define our occlusion aware patch-based virtual view consistency loss as follows:
$L_{\text{virtual}}=L_{\text{SSIM}}(\{\mathbf{c}_{p^{\ast}}\},\{\hat{\mathbf{c}}_{p}\},\{M_{p\rightarrow p^{\ast}}\})+L_{\text{NCC}}(\{\mathbf{c}_{p^{\ast}}\},\{\hat{\mathbf{c}}_{p}\},\{M_{p\rightarrow p^{\ast}}\})$ (6)
Refer to caption Refer to caption
(a) Virtual Patch Sampling (b) Occlusion
Figure 3: Virtual View Patch sampling and occlusion visualization. (a) We first sample a virtual view near the training view. Then, we render a patch from the training view to estimate the depth (Red line), which is projected to the virtual view (Green Line). Finally, we render the virtual patch (Blue line) and compare the rendered RGB to the ground truth RGB in the training patch. (b) A pixel is marked as occluded and excluded from the NCC and SSIM losses, if the projection of its rendered virtual view depth is inconsistent with the training view depth based on the angle difference.
[…]
@@
$\theta_{p^{\ast}\rightarrow p}=\arccos(\mathbf{v}_{p^{\ast}\rightarrow p}\cdot\mathbf{v}_{p}),\quad M_{p\rightarrow p^{\ast}}=\begin{cases}1&\text{if }\theta_{p^{\ast}\rightarrow p}\leq\theta_{\text{thresh}}\\ 0&\text{otherwise}\end{cases}$ (7)
$where\quad\mathbf{v}_{p^{\ast}\rightarrow p}=(\mathbf{X}_{p^{\ast}}-\mathbf{o}_{p})/\|(\mathbf{X}_{p^{\ast}}-\mathbf{o}_{p})\|,\quad\mathbf{X}_{p^{\ast}}=\mathbf{o}_{p^{\ast}}+{d}_{p^{\ast}}\mathbf{v}_{p^{\ast}}$
[…]
@@ 3.4 Density Restriction by Empty Space Pruning with Sparse SfM Geometry
Monocular depth provides useful relative distance information, and SfM points provide metric depth that can be utilized for aligning the monocular depth with 3D space. For each view, we use RANSAC to solve a scale and shift for monocular depth with the projected sparse points to minimize the influence of noise in the points. We then constrain the grid density distribution to a specified interval surrounding the optimal monocular depth maps for each view. This hard restriction effectively prunes out empty space, thereby eliminating the floaters and improving the overall geometry estimation.
Figure 4: Visualization of density restrictions. We show on the left the point cloud reconstruction of our model trained under density restrictions. On the right, we capture a vertical slice of the scene to show how density restrictions improve the geometry. The blue area indicates density-restricted region, and colored area represents the point cloud reconstruction of our method trained under density restrictions. Green and red regions indicate the point cloud of our model trained without density restrictions, where green indicates that the region is inside the density-restricted area, and red indicates that the region is outside the density-restricted area.
@@ 3.5 Training
We start by estimating monocular geometric cues with images using the pretrained Omnidata [5] model. With the sparse points from SfM, we use RANSAC to search for the optimal shift and scale for the monocular depth. We initialize the density restriction by voxelizing the space and labeling each center of a voxel within 20% of the projected monocular depth map, and then exclude sampling outside of labeled voxels. During training, we iterate through images, sample a bunch of patches per reference image, sample one virtual patch for each training patch, and evaluate the loss terms for all patches. We use NerfAcc [15] with modified QFF [14] as our base model for faster training and inference without loss of accuracy, and train with the unified loss $L=\sum\lambda_{:}L_{:}$, with $\lambda_{huber},\lambda_{depth},\lambda_{\nabla depth},\lambda_{normal},\lambda_{\nabla normal},\lambda_{SSIM},\lambda_{NCC}$ being $1.0$, $0.05$, $0.025$, $31E-$ 3, $45E-$ 4, $41E-$ 4, $41E-$ 4. Please see our supplementary material for more details on parameters, model architecture, and training and inference time.
@@ 4 Experiments
Our experiments investigate: (1) the geometric accuracy of existing NeRF methods and our method on the challenging MVS benchmarks, as measured by point cloud metrics and view extrapolation; (2) how each of our contributions and system components affect performance.
We use the training scenes from the ETH3D High-Resolution dataset (ETH3D)  [29] and selected scenes from TanksAndTemples (TnT) [11] as challenging large-scale natural scenes. ETH3D consists of 7 large-scale indoor scenes and 6 outdoor scenes, each sparsely capturing about 35 images on average, which is especially challenging for NeRF. In addition, we select large-scale indoor scenes (Church, Meetingroom) and outdoor scenes (Barn, Courthouse), as well as the advanced testing scenes used in the [38] from TnT to validate our method on densely captured scenes.
Evaluation Protocols: We evaluate the methods on novel view synthesis and geometric inference. For novel view synthesis, we train the model with 90% of the images and treat the remaining 10% images as test views for evaluation. We report Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) [39] as evaluation metrics. For geometric inference, we train on all images, and evaluate using the provided 3D geometry evaluation pipeline [29, 11] on point clouds. To generate point clouds, for radiance based method, we render the expected depth map of each training view, and fuse the depth maps with the scheme proposed by Galliani et al [6]; for SDF based methods, we render the mesh with SDF values using marching cube [17], and sample points from the rendered mesh. For TnT dataset, we only report the geometry inference as all views are densely captured.
@@
Mean / In/ Out Mean / In / Out Mean / In / Out Mean / In / Out
RegNeRF [24] 20.90 0.707 0.439 7.3 / 11.1 / 2.8 6.0 / 9.5 / 1.9 6.4 / 10.0 / 2.2 15.5 / 22.4 / 7.4
FreeNeRF [34] 17.24 0.590 0.581 7.5 / 10.4 / 4.1 2.6 / 3.8 / 1.3 3.3 / 4.7 / 1.7 8.5 / 10.8 / 5.7
MonoSDF [38] 18.85 0.679 0.498 25.2 / 26.2 / 24.0 19.3 / 28.5 / 8.5 20.1 / 26.9 / 12.1 41.1 / 45.2 / 36.4
Neuralengelo [16] 19.53 0.696 0.414 3.3 / 3.4 / 3.2 2.1 / 3.4 / 0.6 2.3 / 3.4 / 1.0 7.2 / 8.0 / 6.2
Ours 20.12 0.720 0.379 36.2 / 45.6 / 25.2 24.4 / 29.8 / 18.2 28.8 / 35.6 / 20.9 46.9 / 52.9 / 40.0
Ours (MVS-Depth) 20.48 0.742 0.341 70.2 / 71.6 / 68.4 53.6 / 58.5 / 47.9 60.4 / 64.0 / 56.3 80.7 / 81.7 / 75.3
Gipuma [6] - - - 86.5 / 89.3 / 83.2 24.9 / 24.6 / 25.3 36.4 / 35.8 / 37.1 49.2 / 47.1 / 51.7
COLMAP [28] - - - 91.9 / 95.0 / 88.2 55.1 / 52.9 / 57.7 67.7 / 66.8 / 68.7 80.5 / 78.5 / 82.9
ACMMP [33] - - - 90.6 / 92.4 / 88.6 77.6 / 79.6 / 75.3 83.4 / 85.3 / 81.3 92.0 / 92.2 / 91.9
Qualitative Results: Figure 5 shows the qualitative results on the ETH3D dataset. Compared to RegNeRF [24] and FreeNeRF [34], our approach provides noise-free, more accurate color, depth and normal maps. Compared to MonoSDF [38] and Neuralangelo [16], we produce high-frequency details (e.g, stairs, lamps) on appearance and geometry. Our fused point cloud shows that our NeRF geometry is more consistent and detailed across multiple views compared to the baselines. Figure 6 provides additional novel view rendered images and meshes among different scenes. Our method estimates clear appearance and accurate geometry, including texture-less, semi-transparent, or reflective surfaces (e.g., tables, glass doors, and windows) where RegNeRF [24] and FreeNeRF [34] fail. Our method is robust to scenes with challenging shape and initialization (Relief_2), while Neuralangelo [16] and original MonoSDF [38] hugely suffer from the setup. For better performance, we reduce the radius of initialized sphere surface (bias parameter) for MonoSDF and achieves much better results, but it still fails to extend to the far end of the gallery. See supplementary material for more comparisons and the local minimum of MonoSDF [38] with original parameters.
Comparison on ETH3D Dataset: We report the quantitative comparison against the baselines on Table 1. Our method achieves the best results for all metrics of geometry and novel view synthesis, except for the second best PSNR. RegNeRF [24] achieves better NVS results than other baselines, while predicting inaccurate geometry. MonoSDF [38] outperforms other baselines for geometry metrics, but the novel view rendering is of low quality. Neuralangelo [16] fails to generate high-quality reconstruction due to the sparsity of the views and not using geometric cues, but it achieves better novel view synthesis results than MonoSDF [38], which shows the advantage of using numerical gradients for hash grid to compute higher-order derivatives (surface normals). These results indicate that current approaches have difficulty rendering high quality images while maintaining correct geometry, and our method significantly outperforms RegNeRF on geometry and MonoSDF on image rendering, showing that our method is effective for both purpose. For a better performance on ETH3D, we modify MonoSDF with a reduced bias term by the authors’ suggestion. Due to GPU limitations, we use a smaller batch size for Neuralangelo than author-provided parameters, which may slightly reduce performance. We provide one comparison of meshes from the two parameters in supplementary material.
@@
Table 2: Comparisons on TnT [11]. We report the $F$-score of each method on Tanks and Temples Dataset. ${}^{\ast}$ indicates results from original papers. For ease of comparison against Neuralangelo [16], we exclude scene Church which is not provided by the authors for comparison. NeuralA. refers to Neuralangelo. Advanced scene results are from the official online evaluation site. We mark the best scoring methods with bold.
Scene Ours MonoSDF ${}^{\ast}$ ${}^{\ast}$ Scene Ours ${}^{\ast}$
@@
Integration of MVS Depth: We find the volumetric (NeRF) based methods provide less accurate geometry than the MVS based methods, especially on the sparse-view setup. Hence, we additionally use ACMMP [33] inferred depth (denoted as $\hat{d}_{p}^{mvs}$ for simplicity) to supervise our rendered depth with an additional L1 Loss $L_{\text{mvs}}=|\hat{d}_{p}^{mvs}-{d}_{p}|$ and weight $\lambda_{\text{mvs}}=0.1$. We do not apply scale and shift for $\hat{d}_{p}^{mvs}$ because MVS depth is metric depth. Marked as “Ours (MVS-Depth)” in Table 1, these additional MVS depths significantly boost the geometry and slightly improve the rendering. However, when supervised with MVS depth, the resulting geometry is less accurate than the MVS fused point cloud. We believe this is because NeRF methods have trouble fitting the MVS depth while also minimizing photometric loss, as per-view photometric loss does not enforce correct geometry as well as MVS cross-view photometric consistency. See Figure 7 for qualitative comparison. We note that “Ours” without additional notation refers to our model with monocular cues but not MVS depth for the whole paper.
Comparison on TnT Dataset: In addition to the sparse views setup, we report the quantitative comparison against the baselines [16, 38, 4] on Table 2. Our method slightly underperforms Neuralangelo [16], indicating the strength of Neuralangelo for scenes with denser captures. We note that Neuralangelo initializes with ground truth points for scene scaling (and is sensitive to this initialization). However, in practice, ground truth is not available, so the scaling of our method is set based on automatically obtained SfM sparse points. In the advanced scenes, we significantly outperform MonoSDF on all four scenes, demonstrating our method’s capability on densely sampled large scale scenes.
Ablation Study: We ablate each component of our method in Tables 3 and 4. The tables show that including all of patch-based training, monocular cues, virtual view-based regularization, and density restriction greatly improves the geometry estimation. Table 3 indicates that monocular cues and the density restriction very lightly decrease the novel view synthesis quality, likely because they prevent the model from using erroneous geometry to create some view-dependent effects that it otherwise has trouble modeling. When using monocular supervision without patch-based training, loss for gradient of depth $L_{\nabla\text{depth}}$ and normals $L_{\nabla\text{normal}}$ are not applied as gradients are less accurate among randomly sampled pixels.
Table 3: Ablations on ETH3D [29]. We report evaluations of different combinations of our components on the ETH3D dataset [29]. Patch denotes using patch-based training, Mono. denotes using monocular geometric cues, Virtual denotes using virtual view-based regularization, and Restr. denotes density restriction. We mark the best-performing combinations for each criterion in bold.
@@
Mean / In / Out Mean / In / Out
18.8 0.695 0.397 6.2 / 10.1 / 1.6 14.2 / 21.6 / 5.7
19.6 0.695 0.393 6.7 / 11.4 / 1.3 15.1 / 24.3 / 4.4
18.2 0.618 0.484 10.3 / 15.0 / 4.8 23.1 / 30.1 / 15.0
20.0 0.723 0.388 8.5 / 11.2 / 5.3 18.3 / 20.8 / 15.5
21.4 0.745 0.382 18.3 / 24.7 / 11.0 33.7 / 39.2 / 27.3
20.9 0.742 0.372 23.1 / 29.5 / 15.7 38.7 / 43.8 / 32.7
20.1 0.720 0.379 28.8 / 35.6 / 20.9 46.9 / 52.9 / 40.0
Table 4: Ablations on TnT [11] dataset Training scenes. We show evaluations of different configurations on the selected TnT scenes. We report the $F$-score for each scene and average precision, recall and F-score for all scenes. We mark the best performing configuration in bold for each evaluation.
Patch Mono. Virtual Restr. Indoor Outdoor Mean
Church Meetingroom Barn Courthouse Prec. Rec. F-score
1.6 4.7 16.1 4.5 9.7 6.2 6.7
1.7 7.7 22.9 6.2 10.2 10.4 9.6
10.2 10.6 24.7 16.4 15.2 17.7 15.5
7.8 13.0 35.7 11.9 15.0 22.2 17.1
2.3 13.7 38.0 22.4 18.7 21.4 19.1
20.3 22.0 49.4 38.3 28.4 38.6 32.5
Limitations: Our method is more geometrically accurate than NeRF/SDF based methods for the tested scenes, but still falls short of MVS systems, even with the MVS supervision. Potential directions for improvements include: guided sampling of virtual view patches; joint inference of geometry with single-view predictions; including per-image terms to better handle lighting effects; expanding material models, e.g. with both diffuse and specular terms per point; and incorporating semantic segmentation. Further, geometrically accurate NeRF-based methods are slow compared to MVS systems, which is a barrier to research and practice.

### v2

@@ 3.3 Patch-based Photometric Consistency over Virtual Views
To better train NeRF models given sparse input views, we enforce an occlusion-aware photometric consistency between training patches and randomly sampled virtual patches. Though designed for sparse-view input, we find that the photometric regularization works for general setups, significantly boosting the performance for both sparse-view (Table 3) and dense-view scenes (Table 4).
Figure 3: Virtual view patch sampling and occlusion visualization. (a) We first sample a virtual center $\textbf{o}_{p^{\ast}}$ near the training center $\textbf{o}_{p}$. We then unproject the training patch to $\{\textbf{X}_{p}\}$ with rendered depth, and project $\{\textbf{X}_{p}\}$ to $\textbf{o}_{p^{\ast}}$ for the virtual patch viewing directions. Color of the virtual patch are rendered and compared to the ground truth RGB in the training patch. (b) We unproject the virtual patch to $\{\textbf{X}_{p^{\ast}}\}$ with virtual rendered depth, and mask pixels based on the angle $\{\theta_{p^{\ast}\rightarrow p}\}$ between $\{\textbf{X}_{p}\}$ to $\textbf{o}_{p}$ and $\{\textbf{X}_{p^{\ast}}\}$ to $\textbf{o}_{p}$. For simplicity, the visualization only contains a single pixel $p$.
As visualized in Fig. 3, for each training patch $P$, we randomly sample a corresponding virtual camera center $o_{p^{\ast}}$ within a fixed distance to $o_{p}$ (0.05×0.05$\times$0.05 × the scene width). The correspondence is determined by unprojecting the pixel from the training view into 3D points, denoted as $\mathbf{X}_{p}=\mathbf{o}_{p}+\hat{d}_{p}\mathbf{v}_{p}$, and then back-projecting $\mathbf{X}_{p}$ into the $\mathbf{o}_{p^{\ast}}$ for the corresponding direction $\mathbf{v}_{p^{\ast}}=(\mathbf{X}_{p}-\mathbf{o}_{p^{\ast}})/\|(\mathbf{X}_{p}-\mathbf{o}_{p^{\ast}})\|$. Given $\mathbf{o}_{p^{\ast}}$ and $\mathbf{v}_{p^{\ast}}$, we render the color $\mathbf{\hat{c}}_{p^{\ast}}$ and depth ${\hat{d}}_{p^{\ast}}$. Since viewing the same point from different perspectives can result in varying occlusion patterns, we additionally apply mask $M_{p\rightarrow p^{\ast}}$ to remove any clearly occluded pixels to eliminate the occlusion impact. We compute $M_{p\rightarrow p^{\ast}}$ by comparing an angle threshold $\theta_{\text{thresh}}$ and the angle $\theta_{p^{\ast}\rightarrow p}$ between rays from $\mathbf{o}_{p}$ to $\mathbf{X}_{p}$ and $\mathbf{o}_{p}$ to $\mathbf{X}_{p^{\ast}}$:
@@
$\begin{split}\theta_{p^{\ast}\rightarrow p}&=\arccos(\mathbf{v}_{p^{\ast}\rightarrow p}\cdot\mathbf{v}_{p}),\\ M_{p\rightarrow p^{\ast}}&=\begin{cases}1&\text{if }\theta_{p^{\ast}\rightarrow p}\leq\theta_{\text{thresh}}\\ 0&\text{otherwise}\end{cases}\\ \end{split}$ (6)
where $\mathbf{v}_{p^{\ast}\rightarrow p}=\frac{(\mathbf{X}_{p^{\ast}}-\mathbf{o}_{p})}{\|(\mathbf{X}_{p^{\ast}}-\mathbf{o}_{p})\|}$ and $\mathbf{X}_{p^{\ast}}=\mathbf{o}_{p^{\ast}}+\hat{d}_{p^{\ast}}\mathbf{v}_{p^{\ast}}$.
The occlusion-aware patch-based consistency loss is defined as:
$\begin{split}L_{\text{virtual}}&=L_{\text{SSIM}}(\{\mathbf{\hat{c}}_{p^{\ast}}\},\{{\mathbf{c}}_{p}\},\{M_{p\rightarrow p^{\ast}}\})\\ &+L_{\text{NCC}}(\{\mathbf{\hat{c}}_{p^{\ast}}\},\{{\mathbf{c}}_{p}\},\{M_{p\rightarrow p^{\ast}}\})\end{split}$ (7)
where $L_{\text{SSIM}}$ and $L_{\text{NCC}}$ measure the structural similarity and normalized cross-correlation between two masked patches respectively. Both losses are robust to view-dependent effects like illumination change. Though we rely on rendered depth for the loss, the occlusion mask will inhibit the loss from incorrectly guiding the optimization in the beginning.
@@ 3.4 Density Restriction by Empty Space Pruning with Sparse SfM Geometry
Monocular depth provides useful relative distance information, and SfM points provide metric depth that can be utilized for aligning the monocular depth with 3D space. For each view, we use RANSAC to solve a scale and shift for monocular depth with the projected sparse points to minimize the influence of noise in the points. We then constrain the grid density distribution to a specified interval around the estimated depth along each ray. This hard restriction effectively prunes out empty space, thereby eliminating the floaters and improving the overall geometry estimation.
Figure 4: Visualization of density restrictions. On the left, we present the point cloud reconstruction of our model trained with density restrictions. On the right, a vertical slice of the reconstructed scene is shown, both with and without density restrictions. The original scene points and color points (green and red) represent our reconstructed point cloud with and without density restrictions, respectively. The blue area denotes density-restricted voxels. With density restrictions, the ground is accurately reconstructed as a plane, whereas without density restrictions, the ground sinks down.
@@ 3.5 Training
We start by estimating monocular geometric cues with images using the pretrained Omnidata [6] model. With the sparse points from SfM, we use RANSAC to find the optimal shift and scale for the monocular depth for density restriction. To initialize the density restriction, we voxelize the space, project the center of each voxel to all training views, label voxels with centers lying within 20% of any monocular depth map, and exclude sampling outside of labeled voxels. During training, we sample $128$ patches per iteration, sample one virtual patch for each training patch, and evaluate the loss terms for all patches. Angle threshold $\delta_{\text{thresh}}$ is set as 10 degrees when estimating occlusion masks. We use NerfAcc [19] with modified QFF [18] as our base model for faster training and inference without loss of accuracy, and train with the unified loss $L=\sum\lambda_{:}L_{:}$, with $\lambda_{\text{rgb}}=1.0$, $\lambda_{\text{depth}}=0.05$, $\lambda_{\nabla\text{depth}}=0.025$, $\lambda_{\text{normal}}=1\times 10^{-3}$, $\lambda_{\nabla\text{normal}}=5\times 10^{-4}$, $\lambda_{\text{SSIM}}=1\times 10^{-4}$, $\lambda_{\text{NCC}}=1\times 10^{-4}$. Please see our supplementary material for more details on parameters and model architecture.
@@ 4 Experiments
Our experiments investigate: (1) the geometric accuracy and rendering quality of existing NeRF methods and our method on the challenging MVS benchmarks, as measured by point cloud metrics and novel view synthesis; (2) how each of our contributions and system components affect performance.
Datasets: We experiment with ETH3D [36] and Tanks and Temples (TnT) [15] because they are among the most challenging benchmarks for MVS. We use the training scenes from the ETH3D High-Resolution dataset (ETH3D)  [36], consisting of 7 large-scale indoor scenes and 6 outdoor scenes. These scenes are sparsely captured, with only 35 images on average, which is especially challenging for NeRF. In addition, we experiment with TnT large-scale indoor scenes (Church, Meetingroom) and outdoor scenes (Barn, Courthouse), as well as the advanced testing scenes used in the [55] to validate our method on densely captured scenes.
Evaluation Protocols: We evaluate the methods on novel view synthesis and geometric inference. For novel view synthesis, we train the model with 90% of the images and treat the remaining 10% images as test views for evaluation. We report Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) [57] as evaluation metrics. For geometric inference, we train on all images, and evaluate using the provided 3D geometry evaluation pipeline [36, 15] on point clouds. To generate point clouds, for density-based methods, we render the expected depth map of each training view, and fuse the depth maps with the scheme proposed by Galliani et al [8]; for SDF-based methods, we render the mesh with SDF values using marching cube [21] and sample points from the rendered mesh. For TnT dataset, we only report the geometry inference as all views are densely captured. Since NeRF papers rarely report F1-scores, we train and evaluate scene models for some of the leading methods in density-based and SDF-based NeRF, using author-provided code, advice, and reported numbers where possible.
@@
Mean / In/ Out Mean / In / Out Mean / In / Out Mean / In / Out Training / Inference
RegNeRF [28] 20.90 0.707 0.439 7.3 / 11.1 / 2.8 6.0 / 9.5 / 1.9 6.4 / 10.0 / 2.2 15.5 / 22.4 / 7.4 14.4 / 36.4 200000
FreeNeRF [48] 17.24 0.590 0.581 7.5 / 10.4 / 4.1 2.6 / 3.8 / 1.3 3.3 / 4.7 / 1.7 8.5 / 10.8 / 5.7 7.0 / 30.4 200000
MonoSDF [55] 18.85 0.679 0.498 25.2 / 26.2 / 24.0 19.3 / 28.5 / 8.5 20.1 / 26.9 / 12.1 41.1 / 45.2 / 36.4 20.9 / 136.0 200000
Neuralangelo [20] 19.53 0.696 0.414 3.3 / 3.4 / 3.2 2.1 / 3.4 / 0.6 2.3 / 3.4 / 1.0 7.2 / 8.0 / 6.2 19.9 / 61.9 200000
Ours 20.12 0.720 0.379 36.2 / 45.6 / 25.2 24.4 / 29.8 / 18.2 28.8 / 35.6 / 20.9 46.9 / 52.9 / 40.0 2.2 / 4.6 50000
Ours (MVS-Depth) 20.48 0.742 0.341 70.2 / 71.6 / 68.4 53.6 / 58.5 / 47.9 60.4 / 64.0 / 56.3 80.7 / 81.7 / 75.3 2.7 / 4.6 50000
Gipuma [8] - - - 86.5 / 89.3 / 83.2 24.9 / 24.6 / 25.3 36.4 / 35.8 / 37.1 49.2 / 47.1 / 51.7 - / - -
COLMAP [34] - - - 91.9 / 95.0 / 88.2 55.1 / 52.9 / 57.7 67.7 / 66.8 / 68.7 80.5 / 78.5 / 82.9 - / - -
ACMMP [47] - - - 90.6 / 92.4 / 88.6 77.6 / 79.6 / 75.3 83.4 / 85.3 / 81.3 92.0 / 92.2 / 91.9 - / - -
4.1 Main Results
Table 1 compares our approach to density-based and SDF-based NeRF approaches and to MVS for both novel view synthesis and point cloud accuracy metrics. Fig. 5 compares RGB, depth, and normal renderings of the NeRF-based approaches, and Fig. 6 shows more examples of meshes and novel view synthesis. Fig. 7 compares generated point clouds to MVS. Table 2 compares geometric accuracy to SDF-based methods on the TnT dataset. More results like novel view video comparisons and sparse-view TnT comparisons are in the supplemental.
Comparison to regularized density-based NeRF: Our approach provides more accurate and detailed novel view synthesis, depth, normals, and meshes than RegNeRF [28] and FreeNeRF [48] in the ETH3D experiments (Tab. 1, Fig. 5, Fig. 6). Our approach particularly outperforms on texture-less, semi-transparent, or reflective surfaces (e.g., tables, glass doors, and windows). RegNeRF is second best and is closest to our approach. While RegNeRF samples patches in virtual views to regularize based on their color and geometry likelihood, our key differentiation is encouraging patch-based geometric consistency with monocular cues and patch-based photometric consistency with virtual views. Our ablation (Table 3) confirms that without using patches or without incorporating both of these types of consistencies, our method achieves F-scores more similar to these others.
@@ 4.1 Main Results
Table 2: Comparisons on TnT [15]. We report the $F$-score of each method on Tanks and Temples Dataset. indicates results from Neuralangelo [20], acquired with scale initialization from ground truth points. ${}^{\text{\textdagger}}$ indicates results from MonoSDF [55]. For ease of comparison against Neuralangelo, we exclude scene Church which is not provided by the authors for comparison. NeuralA. refers to Neuralangelo. Advanced scene results are from the official online evaluation site. We mark the best scoring methods with bold.
Scene Ours MonoSDF NeuralA. NeuralWarp Scene Ours ${}^{\text{\textdagger}}$
@@
Comparison to SDF-based approaches: SDF-based approaches like MonoSDF [55], Neuralangelo [20], and NeuralWarp [5] are well-regularized by solving directly for an implicit surface, and more easily incorporate monocular geometry estimates. SDF-based approaches work very well in some scenes, but have drawbacks of higher dependence on initialization, compute-heavy optimization, and overly smooth surfaces in the results. Qualitative results in the ETH3D scenes (Fig. 5) show that our method better captures details, such as the stairs, lamppost, and tables. In Tab. 1, our method outperforms MonoSDF and Neuralangelo in all novel view synthesis and geometry metrics. Interestingly, Neuralangelo outperforms MonoSDF in NVS but greatly underperforms in geometric accuracy because the sparse views do not provide sufficient regularization to constrain geometry.
The denser views and often simpler scenes make TnT scenes more amenable to SDF-based methods (Tab. 2). In the simpler Training scenes, our method performs best for “Courthouse”, while Neuralangelo performs best for “Meetingroom” and “Barn” and MonoSDF slightly outperforms ours in “Church”. For Advanced TnT scenes, our method consistently outperforms MonoSDF.
Table 3: Ablations on ETH3D [36]. We report evaluations and training time (seconds per 1000 steps) of different combinations of our components on the ETH3D dataset [36]. Patch denotes using patch-based training, Mono. denotes using monocular geometric cues, Virtual denotes using virtual view-based regularization, and Restr. denotes density restriction. We mark the best-performing combinations for each criterion in bold.
@@
18.8 0.695 0.397 6.2 14.2 129
19.6 0.695 0.393 6.7 15.1 150
18.2 0.618 0.484 10.3 23.1 97
20.0 0.723 0.388 11.7 24.0 130
21.4 0.745 0.382 18.3 33.7 165
20.9 0.742 0.372 23.1 38.7 233
20.1 0.720 0.379 28.8 46.9 153
Robustness and compute can also be a concern for SDF-based methods. For example, Neuralangelo requires scale initialization from ground truth point clouds. For MonoSDF, following author advice, we find that reducing the bias parameter (initialized sphere radius) from defaults gives much better results. With this setting, MonoSDF captures the columns in Relief_2 (Fig. 6) but still does not extend to the full depth of the gallery. Neuralangelo is memory-intensive; we had to reduce the batch size from default to train using one A40 GPU (48GB). By contrast, our method uses the same parameters for all experiments, and scale is set automatically based on SfM sparse points, which are typically available for posed images. Our memory and compute requirements are lower. See Tab. 1 and supplemental for details.
Comparison to MVS: MVS methods, such as ACMMP [47], produce very accurate geometry in some portions of scenes but also tend to produce noisy points and incomplete surfaces, and they cannot synthesize realistic novel views. Quantitatively, our method underperforms MVS according to point cloud metrics (Table 1), but qualitatively (Fig. 7) our method’s point cloud is more complete and similarly accurate, especially when using guidance from MVS depth maps (described in Sec. 4.2). Much of the gap is likely due to limitations in extracting the 3D point clouds from the NeRF model. We use a simple approach of rendering expected depth in each view and applying standard fusion techniques as a post-process, while MVS methods often include optimization steps to improve geometric consistency across views. Another possible cause is that multiscale MVS methods better exploit high resolution images.
Table 4: Ablations on TnT [15] dataset Training scenes. We show evaluations of different configurations on the selected TnT scenes. We report the $F$-score for each scene and mark the best performing configuration in bold.
Patch Mono. Virtual Restr. Church Meetingroom Barn Courthouse Mean
1.6 4.7 16.1 4.5 6.7
1.7 7.7 22.9 6.2 9.6
10.2 10.6 24.7 16.4 15.5
7.8 13.0 35.7 11.9 17.1
2.3 13.7 38.0 22.4 19.1
20.3 22.0 49.4 38.3 32.5
4.2 Ablation Study
Key Contributions: We ablate each key contribution of our approach in Tables 3 and 4. Patch-based objectives, monocular cues, virtual view-based regularization, and density restriction are all important for geometry estimation. Table 3 indicates that monocular cues and the density restriction very slightly decrease the novel view synthesis quality, likely because they prevent the model from using erroneous geometry to create some view-dependent effects that it otherwise has trouble modeling. When using monocular supervision without patch-based training, loss for gradient of depth $L_{\nabla\text{depth}}$ and normals $L_{\nabla\text{normal}}$ are not applied as gradients are less accurate among randomly sampled pixels.
Incorporating MVS Depth: We investigate how improved depth maps can affect results of our method. We use ACMMP [47] inferred depth (denoted as $\hat{d}_{p}^{mvs}$ for simplicity) to supervise our rendered depth with an additional L1 Loss $L_{\text{mvs}}=|\hat{d}_{p}^{mvs}-{d}_{p}|$ and weight $\lambda_{\text{mvs}}=0.1$. We do not apply scale and shift for $\hat{d}_{p}^{mvs}$ because MVS depth is metric depth. In Table 1, “Ours (MVS-Depth)” shows that these losses based on MVS depth significantly boost the geometry and slightly improve the rendering.